OpenStudy (anonymous):

find the solution set for this system of equations. x + 3y = 11 AND 2x - 2y =6

OpenStudy (anonymous):

x + 3y = 11 AND 2x - 2y =6 x = 11 - 3y 2(11-3y) - 2y = 6 22 - 6y - 2y = 6 22 - 8y = 6 -8y = 6 -22 -8y = - 16 y = 2 x = 11 -3y x = 11 - 3(2) x = 11 -6 x = 5